OpenStudy (anonymous):

The man who jumps from the bridge falls 5[s] before impact. Determine a) his speed at the time of impact b) his acceleration if the impact lasted only 0.1 [s] c)height of the bridge

OpenStudy (anonymous):

\[v _{f}=v _{i} + at=0+9.8m/s^{2}(5s)=49m/s\]\[a=\Delta v/\Delta t=\frac {0-49m/s} {0.1s}=-490m/s^{2}\]\[h = \frac {(v _{f}^{2} - v _{i}^{2})} {2a}=\frac {(49m/s)^{2}} {19.6m/s^{2}}=122.5m\]